Using arc method calculate the elasticity of demand for oranges when the price

rises from $2 to $3 Price per orange, its demand reduces from 80 thousands oranges

to 70 thousands oranges. Interpret your answer in term of the farmer’s revenue.


1
Expert's answer
2022-05-16T16:08:43-0400

"EOD=\frac{2-3}{2+3}\cdot\frac{80+70}{80-70}=-3."